Mike is on a hunting trip but dosent like to shoot at animals so he places a baseball in a tree right above him. Than shoots a bullet (0.02 kg) at the baseball (0.15 kg). The two objects embed and Rose to a max height of 13.5 meters before coming back down. What was the initial velocity of the bullet?


1
Expert's answer
2021-05-17T10:00:01-0400

"vm_1=u(m_2+m_2),"

"h=\frac{u^2}{2g}\implies u=\sqrt{2gh},"

"v=\frac{\sqrt{2gh}(m_1+m_2)}{m_1}=138~\frac ms."